OpenStudy (anonymous):

example of an equation in two variables with 0 solutions.

OpenStudy (anonymous):

PLEASE HELP

OpenStudy (dan815):

um

OpenStudy (dan815):

2x+y=4x+2y+6

OpenStudy (dan815):

?

OpenStudy (anonymous):

Haaha i really don't even know but thank yo, can you help me out with another one

OpenStudy (dan815):

wait

OpenStudy (dan815):

ill help u -.-

OpenStudy (dan815):

u know equation of lines

OpenStudy (dan815):

y=2x y=2x+1 lets say these 2 equations of lines

OpenStudy (anonymous):

okay

OpenStudy (dan815):

can u see the slopes of those 2 liens are the same

OpenStudy (anonymous):

yea ?

OpenStudy (dan815):

okay these 2 lines have same slope but different y intercept so they will not intersect, which means no solution

OpenStudy (dan815):

you can form an equation with these 2 equations put together

OpenStudy (anonymous):

ohhhhhhhh thats makes sense now

OpenStudy (anonymous):

so how can i find an equation that has two variable but only one answer >

OpenStudy (dan815):

y=2x y=2x+1 --------- y-2x=0 y-2x+1=0 ----> multiply this by 2 just for kicks y-2x=0 2y-4x+2 =0 now equation the 2 zero expressions and we get y-2x=2y-4x+2 here have have an equation which we know cannot be satisfied

OpenStudy (dan815):

okay so similiar to this, now u want to take any 2 lines with different slopes then u know these 2 lines will intersect

OpenStudy (dan815):

y=2x y=3x but lets rewrite this so we sustain the 2 variables Y-2x=0 y-3x=0 now equation the 0 expressions y-2x=y-3x

OpenStudy (dan815):

equate*

OpenStudy (anonymous):

i dont get it ? @dan815

OpenStudy (dan815):

okay u know 2 lines with different slopes always meet

OpenStudy (dan815):



OpenStudy (anonymous):

yea

OpenStudy (dan815):

so take any 2 line equations with different slope

OpenStudy (anonymous):

and theyll have the same answer ?

OpenStudy (dan815):

yeah but re arragen it so u still have the 2 variables

OpenStudy (dan815):

i was thinking uhh.. what grade are u in... have u learnt parabolas

OpenStudy (anonymous):

im so confused

OpenStudy (dan815):

because u can do this with parabolas too

OpenStudy (dan815):